OpenStudy (anonymous):

If there is a 10% chance of sun tomorrow and a 20% chance of wind and no sun, what is the probability that it is windy, given that it is not sunny? Round your answer to the nearest percent.

OpenStudy (anonymous):

help please

OpenStudy (kropot72):

Probability of sun P(S) = 0.1 Therefore the probability of no sun P(NS) = 1 - 0.1 = 0.9 Let the probability of wind, given that it is not sunny, be P(W). It is given that \[\large P(W \cap NS)=0.2\] However \[\large P(W \cap NS)=P(W) \times P(NS)\] Therefore we can write \[\large P(W)=\frac{P(W\cap NS)}{P(NS)}=\frac{0.2}{0.9}=you\ can\ calculate\] When you have calculated the value of P(W) you need to multiply the decimal value by 100 to convert it to a percentage.